OpenStudy (anonymous):

Convert 846 km/h to miles per hour .a.770 mi/h. c. 525 mi/h b.684 mi/h d.465 mi/h

OpenStudy (austinl):

\(\displaystyle \frac{846km}{1h}\times\frac{1mile}{1.60934km}\) This means that the km will be left with miles/hour. Just multiply top and bottom and divide and you should get your answer.

OpenStudy (anonymous):

Thanks.

OpenStudy (austinl):

No problem.